OpenStudy (chris215):

given G(x)= integral 4^x sqrt(1+t^2) dt find G '(x)

OpenStudy (chris215):

\[G(x)= \int\limits_{4}^{x}\sqrt{1+t^2} dt\]

OpenStudy (chris215):

I got \[\sqrt{1+t^2}\]